On visiting the town where I was born, I realised that I had been away for one-third of my life. How long will it be before I will have been away for half of my life?: On visiting the town where I was born, I realised that I had been away for one-third of my life. How long will it be before I will have been away for half of my life?
Answer by stanbon(18015)

1/2 - 1/3 = 3/6 - 2/6 = 1/6 of your current age
=================
Cheers,
Stan H.